# HopCast: Calibration of Autoregressive Dynamics Models

### v4

@@ 4.2 Training
The correction model ($M_{x}$) consists of an Encoder ($m_{x}$) and an MHN ($\text{MHN}_{x}$). The Encoder $m_{x}$ takes in as input the context state, such as $c_{4}^{3}$, and outputs a $d$-dimensional embedding. We choose $d=4$ in our experiments, though other values of $d$ are expected to perform comparably [Appendix D.4]. We use the same $m_{x}$ to construct embedded queries $\mathbf{Q}_{\psi}$ and keys $\mathbf{K}_{\psi}$ from $\mathbf{Q}$ and $\mathbf{K}$, respectively. The $\text{MHN}_{x}$ uses an attention mechanism (Vaswani, 2017) to construct an association matrix ($\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$) based on the similarity of elements in $\mathbf{Q}_{\psi}$ with elements in $\mathbf{K}_{\psi}$. Mathematically,
@@
The association matrix ($\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$) is shown in the Training block of Fig. 2 with $\mathtt{S_{L}}=5$. We mask the association from Autoregressive context state to its Integrated context state. The $(i,j)^{th}$ entry of $\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$ shows the similarity of $i^{th}$ element of $\mathbf{Q}_{\psi}$ with $j^{th}$ element of $\mathbf{K}_{\psi}$. We use $\mathtt{softmax}$ over $\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}[i,:]$ to get association weights for $i^{th}$ element of $\mathbf{Q}_{\psi}$. Therefore, each row of $\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$ sums up to one. The loss function $\mathcal{L}$ is used to learn the parameters of the $m_{x}$ via backpropagation, where the $\hat{\mathbf{V}}_{x}$ denotes the predicted errors.
@@ 4.3 Inference
At inference time, a set of $K$ keys ($\mathbf{K}_{mem}$) with the corresponding values ($\mathbf{V}_{mem}^{x}$) is sampled from $\mathcal{S}_{K}^{\text{Train}}$ and $\mathcal{S}_{V_{x}}^{\text{Train}}$, respectively. For instance, a randomly sampled set of keys and values would be $\mathbf{K}_{mem}=(c_{4}^{3},c_{5}^{1},c_{7}^{5},\cdots,c_{5}^{2})$ and $\mathbf{V}_{mem}^{x}=(e_{x_{4}^{3}},e_{x_{5}^{1}},e_{x_{7}^{5}},\cdots,e_{x_{5}^{2}})$, respectively. The trained $m_{x}$ is used to encode the $\mathbf{K}_{mem}$ as $\mathbf{K}_{mem}^{\psi}$. The $\mathbf{K}_{mem}^{\psi}$ and $V_{mem}^{x}$ constitute the Association memory of MHN. It is shown in the Inference block in Fig. 2. We loaded 2000 keys in memory for evaluation [Appendix D.3]. Once the memory is set up, the correction model ($M_{x}$) takes in the context state $(\bar{x}_{1},\bar{y}_{1},\bar{x}_{t},\bar{y}_{t},t)$ as a query $\mathbf{Q}_{t}$ at $t^{th}$ time step during autoregression and encodes it as $\mathbf{Q}_{t}^{\psi}$. The association weights ($\mathbf{a}_{x}^{t}$) at $t^{th}$ timestep are,
@@
where $\mathbf{a}_{x}^{t}=\{p_{i}\}_{i=1}^{K}$, $\sum_{i=1}^{K}p_{i}=1$, and $K$ denotes the number of keys in memory. A set of $s$ categories is sampled as: $\{b_{i}\}_{i=1}^{s}\sim\text{Categorical}(\mathbf{a}_{x}^{t})$. A set of errors ($E_{\bar{x}}$) is selected from $\mathbf{V}_{mem}^{x}$ based on sampled categories $\{b_{i}\}_{i=1}^{s}$, i.e., $E_{\bar{x}}=\{V_{mem}^{x}[b_{i}]\}_{i=1}^{s}$. A sampled set of errors $E_{\bar{x}}$ is shown in Fig. 3 for the x-output of the Predictor $B$ for the Lorenz system.
@@ 5.1 Evaluation Metrics
Calibration Error (Kuleshov et al., 2018) (CE). A calibrated model has observed fractions ($\hat{p}$) of its predicted random variable match with the expected fractions ($p$). The difference between these two is calibration error. To evaluate CE, $m(=9)$ equally spaced prediction intervals are chosen from 10% to 90%. For each prediction interval, we count the number of times the observed variable falls between intervals. Mathematically, $\text{CE}=\sum_{i=1}^{m}(\hat{p}_{i}-p_{i})^{2}$.
Prediction Interval Width (Auer et al., 2023) (PI-Width). A calibrated model should also be sharp. Sometimes, it is trivial to reduce the CE by always predicting the expected value of a random variable (Kuleshov et al., 2018). A good predictor should be calibrated and sharp. Hence, we propose to use PI-Width as a measure of sharpness. Mathematically,
$\text{PI-Width}=\frac{1}{m}\frac{1}{v}\sum_{i=1}^{m}\sum_{j=1}^{v}|U_{j}^{i}-L_{j}^{i}|,$ (9)
where $m=9$ is for equally spaced PI, and $v$ denotes the number of validation data points; i.e., $v=|\mathcal{S}_{Q}^{\text{Inference}}|$.
@@ 5.2 Datasets
Table 1: Details about data generation
Model $\Delta t$ Timesteps Trajectories
Lotka Volterra 0.1 300 500
Lorenz 0.01 300 1000
FHN 0.5 400 350
Lorenz95 0.01 300 666
Glycolytic 0.01 400 750
@@ 5.3 Baselines
The baseline methods include three uncertainty propagation approaches, i.e., Expectation, Moment Matching, and Trajectory Sampling. These approaches are used to propagate uncertainty with Probabilistic Ensembles (Chua et al., 2018). The details of Probabilistic Ensembles’ training and uncertainty propagation methods are presented in Appendix A.
@@ 6 Results
Model HopCast Probabilistic Ensemble Expectation Moment Matching Trajectory Sampling Metrics MSE PI-Width CE MSE PI-Width CE MSE PI-Width CE MSE PI-Width CE Lotka
@@
HopCast lets us control the confidence of the model with precise control over the width of PI based on a concept we introduce called Attention Span. The Attention Span can be increased/decreased by increasing/decreasing the $\mathtt{S_{L}}$ of Q and K in $\textbf{A}_{\mathtt{S_{L}}\times\mathtt{S_{L}}}$. To demonstrate that, we collect a dataset $H=\{(x_{i},y_{i})\}_{i=1}^{6000}$ using the sine function with heteroscedastic noise (Chua et al., 2018) shown in equation 10 and Fig. 4(a). The $x_{i}$’s are uniformly sampled from the following domain $\frac{5\pi}{2},-\pi]\cup[\pi,\frac{5\pi}{2}]$.
@@ 7 Hyperparameter Tuning
$\mathtt{S_{L}}$ (Sequence Length) is a hyperparameter that must be tuned for calibrated uncertainty. We propose Algorithm 1 to hand-tune it based on the intuition that the uncertainty increases as we increase $\mathtt{S_{L}}$ as discussed in section 6 and shown in Fig. 7. We pick a relatively smaller (e.g., 10) and larger (e.g., 1000) $\mathtt{S_{L}}$ and expect the model to show overconfidence and underconfidence on the calibration curve, respectively. Next, we expect three cases. First, we pick $\mathtt{S_{L}}=500$, which is the optimal $\mathtt{S_{L}}$ because it is very close to perfect calibration (dark dotted line in Fig. 7). Second, our choice leads to overconfidence, e.g., $\mathtt{S_{L}}=400$. Third, there is an underconfident model, e.g., $\mathtt{S_{L}}=600$. In the first case, we have the calibrated uncertainty. In the second case, an $\mathtt{S_{L}}$ slightly bigger than 400 would be a good choice. In the third case, an $\mathtt{S_{L}}$ relatively smaller than 600 is a good choice. If it is the second or third case, the $\mathtt{S_{L}}$ needs to be re-tuned and we expect three cases again. We iteratively do this with the help of the calibration curve to find optimal $\mathtt{S_{L}}$ within or less than five to six trials. Appendix D.7 has additional results highlighting the impact of $\mathtt{S_{L}}$ on uncertainty.
Figure 7: Calibration curves of y-output of Lorenz at different $\mathtt{S_{L}}$. The dark dotted line corresponds to perfect calibration.
The number of models in an ensemble is a hyperparameter. As discussed in the section 6, the ensembles tend to build up the uncertainty until it plateaus [Appendix D.6]. We keep adding models to the ensemble until it provides calibrated uncertainty. In some cases, the uncertainty saturates while the model is overconfident. In those cases, the number of optimal models is very high, e.g., 8 or 7. Appendix E contains additional details about hyperparameters.
@@ 8 Conclusions
We proposed a Predictor-Corrector mechanism for autoregressive dynamics models to correct the Predictors’ predictions and generate calibrated prediction intervals for it at any timestep during autoregression. HopCast performs competitively well against the alternative approach (i.e., uncertainty propagation), giving accurate and calibrated autoregressive dynamics models. Out of three uncertainty propagation approaches, i.e., Trajectory Sampling, Moment Matching, and Expectation, Trajectory Sampling performs competitively well against ours. This is due to its nature of taking predicted uncertainty into account during propagation. HopCast offers a lower calibration and prediction error with sharper prediction intervals. The sharper prediction intervals result directly from modeling context-specific errors, and the lower prediction error is the consequence of modeling errors in the form of a Corrector. In addition, we introduce a concept called Attention Span that gives precise control over the width of prediction intervals with a hyperparameter we introduce called Sequence Length ($\mathtt{S_{L}}$).
@@
One of the limitations of our work is due to the use of explicit timestep IDs (i.e., $t$ =1,2,3, etc.) in the context state ($c_{t}^{n}$). At inference time, the Corrector will only generate calibrated prediction intervals for a trajectory length equal to or less than what it was trained on. Moreover, the Corrector was trained on trajectories sampled at a fixed time interval $\Delta t$ and does not generalize to other sampling intervals. A modified version of the Corrector could be made discretization-invariant. We hand-tuned the $\mathtt{S_{L}}$ hyperparameter, training multiple models to get a lower calibration error. An alternative way to get calibrated uncertainty with HopCast without tuning $\mathtt{S_{L}}$ hyperparameter is an interesting future work.
@@ Appendix A Baseline Methods
To construct Probabilistic Ensembles, we train a population of fully connected feedforward models, i.e., $\Omega=\{f_{i}\}_{i}$, where $f_{i}$ represents the $i^{th}$ model within the population. Each model is trained with Gaussian negative log-likelihood (NLL) loss as demonstrated in Lakshminarayanan et al. (2017). Each model predicts a multivariate normal (MVN) over the next state $\mathbf{s}_{t+1}$ given the previous state of the system $\mathbf{s}_{t}$. Formally, $\mathbf{s}_{t+1}\sim\text{MVN}(\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})$. For the LV system, $\bm{\mu}_{t+1}=(\mu_{t+1}^{x},\mu_{t+1}^{y})$ and $\bm{\Sigma}_{t+1}=\text{diag}(\sigma_{t+1}^{x},\sigma_{t+1}^{y})$.
@@
At inference time, a sample of $M$ models $\{f_{m}\}_{m=1}^{M}$ can be taken from the population ($\Omega$) to form a Probabilistic Ensemble. All three propagation approaches use a particle-based approach, where the core idea is to generate multiple trajectories via autoregression starting from the same initial condition (Chua et al., 2018). Let $\mathbf{s}_{t}^{p,m}$ denote the state of the system associated with $p^{th}$ particle from $m^{th}$ model at $t^{th}$ time step. Each model $f_{m}$ receives $P$ copies of the same initial condition $\{\mathbf{s}_{t=0}^{p,m}\}_{p=1}^{P}$, where $P$ denotes the number of particles assigned to model $f_{m}$. The propagation of uncertainty starts with PMP*M copies of the same initial conditions across $\{f_{m}\}_{m=1}^{M}$ models and results in PMP*M number of trajectories. The diversity in these trajectories will come from two sources, i.e., $M$ different models and their probabilistic predictions. These diverse trajectories will be used to derive prediction intervals. In contrast, HopCast used a set of errors (e.g., $E_{\bar{x}}$ & $E_{\bar{y}}$) to derive prediction intervals.

### v5

@@ 4.2 Training
Figure 2 illustrates how the correction model ($M_{x}$) retrieves context-dependent errors using an Encoder ($m_{x}$) and a Modern Hopfield Network ($\text{MHN}_{x}$): (i) the $m_{x}$ maps each context state into a $d$-dimensional embedding, and (ii) the $\text{MHN}_{x}$ forms associations between embedded queries and keys and retrieves the corresponding error values. Specifically, the $m_{x}$ takes in as input the context state, such as $\mathbf{c}_{4}^{3}$, and outputs a $d$-dimensional embedding. We choose $d=4$ in our experiments, though other values of $d$ are expected to perform comparably [Appendix E.6]. We use the same $m_{x}$ to construct embedded queries $\mathbf{Q}_{\psi}$ and keys $\mathbf{K}_{\psi}$ from $\mathbf{Q}$ and $\mathbf{K}$, respectively. These appear as row/column labels in the Training block of Fig. 2. The $\text{MHN}_{x}$ uses an attention mechanism (vaswani2017attention) to construct an association matrix ($\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$) based on the similarity of elements in $\mathbf{Q}_{\psi}$ with elements in $\mathbf{K}_{\psi}$. Mathematically,
@@
The association matrix ($\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$) is shown in the Training block of Fig. 2 with $\mathtt{S_{L}}=5$. We mask the association from Autoregressive context state to its Integrated context state. This masking is reflected in the zeroed diagonal entries of the association matrix. The $(i,j)^{th}$ entry of $\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$ shows the similarity of $i^{th}$ element of $\mathbf{Q}_{\psi}$ with $j^{th}$ element of $\mathbf{K}_{\psi}$. We use $\mathtt{softmax}$ over $\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}[i,:]$ to get association weights for $i^{th}$ element of $\mathbf{Q}_{\psi}$. Therefore, each row of $\mathbf{A}_{{\mathtt{S_{L}}}\times{\mathtt{S_{L}}}}$ sums up to one. The loss function $\mathcal{L}$ is used to learn the parameters of the $m_{x}$ via backpropagation, where the $\hat{\mathbf{V}}_{x}$ denotes the predicted errors.
@@ 4.3 Inference
At inference time, the correction model ($M_{x}$) retrieves past errors based on the similarity between the encoded query and the encoded keys stored in the $\text{MHN}_{x}$ Association memory. To build the Association memory, a set of $K$ keys ($\mathbf{K}_{mem}$) with the corresponding values ($\mathbf{V}_{mem}^{x}$) is sampled from $\mathcal{S}_{K}^{\text{Train}}$ and $\mathcal{S}_{V_{x}}^{\text{Train}}$, respectively. For instance, a randomly sampled set of keys and values would be $\mathbf{K}_{mem}=(\mathbf{c}_{4}^{3},\mathbf{c}_{5}^{1},\mathbf{c}_{7}^{5},\cdots,\mathbf{c}_{5}^{2})$ and $\mathbf{V}_{mem}^{x}=(e_{x_{4}^{3}},e_{x_{5}^{1}},e_{x_{7}^{5}},\cdots,e_{x_{5}^{2}})$, respectively. The trained Encoder $m_{x}$ is used to encode the $\mathbf{K}_{mem}$ as $\mathbf{K}_{mem}^{\psi}$. The $\mathbf{K}_{mem}^{\psi}$ and $V_{mem}^{x}$ constitute the Association memory of $\text{MHN}_{x}$. It is shown in the Inference block in Fig. 2. We loaded 2000 keys in memory for evaluation [Appendix E.5]. Once the memory is set up, the correction model ($M_{x}$) takes in the context state $(\bar{x}_{1},\bar{y}_{1},\bar{x}_{t},\bar{y}_{t},t)$ as a query $\mathbf{Q}_{t}$ at $t^{th}$ time step during autoregression and encodes it as $\mathbf{Q}_{t}^{\psi}$. The association weights ($\mathbf{a}_{x}^{t}$) at $t^{th}$ timestep are,
@@
where $\mathbf{a}_{x}^{t}=(p_{1},\cdots,p_{K})$, $\sum_{i=1}^{K}p_{i}=1$, and $K$ denotes the number of keys in memory. The weights $\mathbf{a}_{x}^{t}$ show how strongly the query $\mathbf{Q}_{t}^{\psi}$ matches each entry in $\mathbf{K}_{mem}^{\psi}$. A set of $s$ categories is sampled as: $\{b_{i}\}_{i=1}^{s}\sim\text{Categorical}(\mathbf{a}_{x}^{t})$. We select $s=1000$ in our experiments [Appendix E.3]. A set of errors ($E_{\bar{x}}$) is selected from $\mathbf{V}_{mem}^{x}$ based on sampled categories $\{b_{i}\}_{i=1}^{s}$, i.e., $E_{\bar{x}}=\{V_{mem}^{x}[b_{i}]\}_{i=1}^{s}$. A sampled set of errors $E_{\bar{x}}$ is shown in Fig. 3 for the x-output of the Predictor $B$ for the Lorenz system. An alternative approach to sampling from $\mathbf{a}_{x}^{t}$ is to retrieve top-$k$ probabilities. We provide an ablation study on this in Appendix E.4.
@@ 5.1 Evaluation Metrics
Calibration Error (kuleshov2018accurate) (CE). A calibrated model has observed fractions ($\hat{p}$) of its predicted random variable match with the expected fractions ($p$). The difference between these two is calibration error. To evaluate CE, $w(=9)$ equally spaced prediction intervals are chosen from 10% to 90%. For each prediction interval, we count the number of times the observed variable falls between intervals. Mathematically, $\text{CE}=\sum_{i=1}^{w}(\hat{p}_{i}-p_{i})^{2}$.
Prediction Interval Width (auer2023conformal) (PI-Width). A calibrated model should also be sharp. Sometimes, it is trivial to reduce the CE by always predicting the expected value of a random variable (kuleshov2018accurate). A good predictor should be calibrated and sharp. Hence, we propose to use PI-Width as a measure of sharpness. Mathematically,
$\text{PI-Width}=\frac{1}{w}\frac{1}{v}\sum_{i=1}^{w}\sum_{j=1}^{v}|U_{j}^{i}-L_{j}^{i}|,$ (11)
where $w=9$ is for equally spaced PI, and $v$ denotes the number of validation data points; i.e., $v=|\mathcal{S}_{Q}^{\text{Inference}}|$.
@@ 5.3 Baselines
The baseline methods include three uncertainty propagation approaches, i.e., Expectation, Moment Matching, and Trajectory Sampling. These approaches are used to propagate uncertainty with Probabilistic Ensembles (chua2018deep). The details of Probabilistic Ensembles’ training and uncertainty propagation methods are presented in Appendix A.
@@ 6 Results
Model HopCast Probabilistic Ensemble (PE) Expectation Moment Matching Trajectory Sampling Metrics MSE PI-Width CE MSE PI-Width CE MSE PI-Width CE MSE PI-Width CE Lotka
@@
HopCast lets us control the confidence of the model with precise control over the width of PI based on a concept we introduce called Attention Span. The Attention Span can be increased/decreased by increasing/decreasing the $\mathtt{S_{L}}$ of Q and K in $\textbf{A}_{\mathtt{S_{L}}\times\mathtt{S_{L}}}$. To demonstrate that, we collect a dataset $H=\{(x_{i},y_{i})\}_{i=1}^{6000}$ using the sine function with heteroscedastic noise (chua2018deep) shown in equation 12 and Fig. 4(a). The $x_{i}$’s are uniformly sampled from the following domain $\frac{5\pi}{2},-\pi]\cup[\pi,\frac{5\pi}{2}]$.
@@ 7 Hyperparameter Tuning
$\mathtt{S_{L}}$ (Sequence Length) is a hyperparameter that must be tuned for calibrated uncertainty. We propose Algorithm 1 to tune it based on the intuition that uncertainty typically increases with larger $\mathtt{S_{L}}$ (Section 6; Fig. 7). A small value (e.g., 10) produces overconfident estimates, while a large value (e.g., 1000) leads to underconfidence. Intermediate values such as $\mathtt{S_{L}}=500$ yield near-perfect calibration, whereas $\mathtt{S_{L}}=400$ and $\mathtt{S_{L}}=600$ show overconfidence and underconfidence, respectively. We iteratively refine $\mathtt{S_{L}}$ with the calibration curve and typically converge to an optimal setting within a few trials. Appendix E.9 provides additional experiments.
The number of models in an ensemble is a hyperparameter. As discussed in the section 6, the ensembles tend to build up the uncertainty until it plateaus [Appendix E.8]. We keep adding models to the ensemble until it provides calibrated uncertainty. In some cases, the uncertainty saturates while the model is overconfident. In those cases, the number of optimal models is very high, e.g., 8 or 7. Appendix F contains additional details about hyperparameters.
Figure 7: Calibration curves of the $y$-output of Lorenz for different $\mathtt{S_{L}}$.
@@
8 Model-Based Reinforcement Learning
We compare the performance of HopCast and Probabilistic Ensembles as the uncertainty-aware dynamics models within a model-based reinforcement learning (RL) algorithm. PETS (Probabilistic Ensembles with Trajectory Sampling) is a widely used model-based RL algorithm that relies on Probabilistic Ensembles to model the transition dynamics and propagate uncertainty for planning via Model Predictive Control (MPC) (chua2018deepreinforcementlearninghandful). In PETS, the planning quality is heavily influenced by the accuracy and calibration of the dynamics model, model bias compounds rapidly during long-horizon rollouts, and miscalibrated dynamics models lead the planner to choose overly optimistic or overly conservative action sequences.
Our earlier experiments in Table 1 demonstrate that HopCast produces more accurate and better-calibrated multi-step predictions than Probabilistic Ensembles, which utilize different uncertainty propagation methods, i.e., Expectation (E) A.1, Moment Matching (MM) A.3, and Trajectory Sampling (TS) A.2. In Table 2, we show the performance of HopCast on four standard control tasks when it is used as a drop-in replacement for the Probabilistic Ensembles in the PETS algorithm. The results for baselines, i.e., Probabilistic Ensembles & Deterministic Ensembles, with three uncertainty propagation methods, i.e., Expectation, Moment Matching, & Trajectory Sampling, are taken from chua2018deepreinforcementlearninghandful  [^2] Code available at: https://github.com/kchua/handful-of-trials. The HopCast yields comparable performance to Probabilistic Ensembles on Cartpole and 7-DOF Reacher, while it outperforms on 7-DOF Pusher and Half-Cheetah tasks. The Probabilistic Ensembles, which utilize uncertainty propagation methods, outperform on Cartpole and 7-DOF Reacher, showing their competitiveness against HopCast. However, the Deterministic Ensembles exhibit competitive performance in a relatively low-dimensional setting, such as Cartpole, but show performance degradation on high-dimensional tasks. These experiments demonstrate the efficacy of our proposed methodology HopCast within a model-based RL algorithm on various control tasks.
Table 2: The performance of different tasks in terms of rewards with different dynamics models, i.e., Probabilistic Ensembles (PE) & Deterministic Ensembles (DE), and various uncertainty propagation methods, i.e., Expectation (E), Moment Matching ($M$), & Trajectory Sampling (TS). The HopCast shows the results with our proposed methodology as a drop-in replacement of PE within the PETS planning algorithm. The best results are shown in bold.
Task Model HopCast PE-E PE-MM PE-TS DE-E DE-MM DE-TS Cartpole $181\pm 2.3$ $180$ $181$ $\mathbf{183}$ $179$ $177$ $181$ 7-DOF Pusher $\mathbf{-45\pm 4.3}$ 48-48 46-46 46-46 95-95 97-97 93-93 7-DOF Reacher $\pm 3.4 44-44$ 45-45 $\mathbf{-43}$ 93-93 94-94 96-96 Half-cheetah $\mathbf{7170\pm 40}$ $5700$ $200$ $7100$ $3800$ $190$ $3950$
@@ 9 Conclusions
We proposed a Predictor-Corrector mechanism for autoregressive dynamics models to correct the Predictors’ predictions and generate calibrated prediction intervals for it at any timestep during autoregression. HopCast performs competitively well against the alternative approach (i.e., uncertainty propagation), giving accurate and calibrated autoregressive dynamics models. Out of three uncertainty propagation approaches, i.e., Trajectory Sampling, Moment Matching, and Expectation, Trajectory Sampling performs competitively well against ours. This is due to its nature of taking predicted uncertainty into account during propagation. HopCast offers a lower calibration and prediction error with sharper prediction intervals. The sharper prediction intervals result directly from modeling context-specific errors, and the lower prediction error is the consequence of modeling errors in the form of a Corrector. In addition, we introduce a concept called Attention Span that gives precise control over the width of prediction intervals with a hyperparameter we introduce called Sequence Length ($\mathtt{S_{L}}$). We also deploy HopCast within a model-based reinforcement learning planner as an alternative to Probabilistic Ensembles, demonstrating its competitive performance on diverse control tasks.
@@
One of the limitations of our work is due to the use of explicit timestep IDs (i.e., $t$ =1,2,3, etc.) in the context state ($c_{t}^{n}$). At inference time, the Corrector will only generate calibrated prediction intervals for a trajectory length equal to or less than what it was trained on. Moreover, the Corrector was trained on trajectories sampled at a fixed time interval $\Delta t$ and does not generalize to other sampling intervals. A modified version of the Corrector could be made discretization-invariant. We hand-tuned the $\mathtt{S_{L}}$ hyperparameter, training multiple models to get a lower calibration error. An alternative approach to obtaining calibrated uncertainty with HopCast without tuning the $\mathtt{S_{L}}$ hyperparameter is an interesting area for future work.
10 Acknowledgements
This work was partially supported by the National Aeronautics and Space Administration, USA, under grant 80NSSC24CA037.
@@ Appendix A Baseline Methods
To construct Probabilistic Ensembles, we train a population of fully connected feedforward models, i.e., $\Omega=\{f_{i}\}_{i}$, where $f_{i}$ represents the $i^{th}$ model within the population. Each model is trained with Gaussian negative log-likelihood (NLL) loss as demonstrated in lakshminarayanan2017simple. Each model predicts a Multivariate Normal (MVN) over the next state $\mathbf{s}_{t+1}$ given the previous state of the system $\mathbf{s}_{t}$. Formally, $\mathbf{s}_{t+1}\sim\text{MVN}(\bm{\mu}_{t+1},\bm{\Sigma}_{t+1})$. For the LV system, $\bm{\mu}_{t+1}=(\mu_{t+1}^{x},\mu_{t+1}^{y})$ and $\bm{\Sigma}_{t+1}=\text{diag}(\sigma_{t+1}^{x},\sigma_{t+1}^{y})$.
@@
At inference time, a sample of $M$ models $\{f_{m}\}_{m=1}^{M}$ can be taken from the population ($\Omega$) to form a Probabilistic Ensemble. All three propagation approaches use a particle-based approach, where the core idea is to generate multiple trajectories via autoregression starting from the same initial condition (chua2018deep). Let $\mathbf{s}_{t}^{p,m}$ denote the state of the system associated with $p^{th}$ particle from $m^{th}$ model at $t^{th}$ time step. Each model $f_{m}$ receives $P$ copies of the same initial condition $\{\mathbf{s}_{t=0}^{p,m}\}_{p=1}^{P}$, where $P$ denotes the number of particles assigned to model $f_{m}$. The propagation of uncertainty starts with PMP*M copies of the same initial conditions across $\{f_{m}\}_{m=1}^{M}$ models and results in PMP*M number of trajectories. The diversity in these trajectories will come from two sources, i.e., $M$ different models and their probabilistic predictions. These diverse trajectories will be used to derive prediction intervals. In contrast, HopCast used a set of errors (e.g., $E_{\bar{x}}$ & $E_{\bar{y}}$) to derive prediction intervals.